solve by elimination method: 3x - 9y = 2 and 3x - 9y = 3
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

3x - 9y = 2
-(3x - 9y = 3) SUBTRACT THESE EQUATIONS.
--------------------
0X+0Y=-1
0=1 THERE IS NO UNIQUE SOLUTION FOR X & Y.

Answer by Alan3354(69443)

There is no solution, unique or otherwise.